OpenStudy (anonymous):

factor the expression 2x^2 + 9x +10

OpenStudy (amistre64):

given the form ax^2 +bx +c; we can see if it factors into integers by: factors of a*c that add up to b 2(10) = 20 1,20; 1+20 = 21 2,10; 2+10 = 12 4,5 ; 4+5 = 9 thatll do it since we multiplied thru by "a"; we have to divide it back out (x +4/2)(x+5/2) ; simplify and if its still a fraction place the bottom in front (x+2) (2x+5)